OpenStudy (anonymous):

Simplify the expression. 22 • 23

OpenStudy (anonymous):



OpenStudy (anonymous):

it didnt right

OpenStudy (anonymous):

\[2^2 * 2^3\] \[2^2 = 2*2\] \[2^3 = 2*2*2\]

jigglypuff314 (jigglypuff314):

the powers add when multiplied like that for example x^4 * x^3 = x^(4+3) = x^7 so 2^2 * 2^3 = 2^(2+3) = 2^__

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

\[2^2\]is saying 2*2 \[2^3\] is saying 2*2*2 what is 2*2? what is 2*2*2?

jigglypuff314 (jigglypuff314):

mmm like x^2 = x times x x^3 = x times x times x so x^2 times x^3 = (x*x)*(x*x*x) = x^5

OpenStudy (anonymous):

\[so 2^{3}\]

jigglypuff314 (jigglypuff314):

mmm 2^2 * 2^3 is not 2^3 :/ try again?

OpenStudy (anonymous):

32

jigglypuff314 (jigglypuff314):

yep, that would be correct :) 32 = 2^5

OpenStudy (anonymous):

yay